# references.bib Discovering Dynamic Symbolic Policies with Genetic Programming

### v2

@@ 2.6.1 Stochastic harmonic oscillator
The stochastic harmonic oscillator (SHO) is a linear model that describes the position and velocity of an oscillating mass in a single dimension under random perturbations. The SHO is visualized in Figure 2A and the dynamics of the SHO are given by
@@
with and . Here, punishes the policy for the distance between and some target state whereas penalizes the control strength. The initial condition is randomly sampled from a normal distribution with zero mean and variances 3 and 1 for the position and velocity respectively. The target position is chosen randomly by sampling uniformly between -3 and 3 in each trial, while the target velocity is always set to zero. As the optimal LQG control is linear, the function set includes basic arithmetic operators.
GP and the baselines are tested on the SHO in three different experiments. In the first experiment the policies interact with the standard setting of the SHO. In this experiment is specified with and and the state observations are noisy. In the second experiment the state of the SHO is partially observed, as the policies only observe the position while the velocity is unknown. remains the same in this experiment. In the third experiment, the policies are evaluated on generalisability. In each trial and are different, specifically and . In this experiment both the position and velocity are observed again. The best dynamic symbolic policy found in the third experiment is also tested on SHO that change over time to analyse how well the latent states of the dynamic policy adapt to these changes. In the experiments with observation noise, the parameters of the readout function are set to and (see Equation 2). In experiments with partial state observability, is adapted such that only the relevant observations are returned and the dimension of is decreased to match the number of observations.
@@ 2.6.2 Acrobot swing up task
The acrobot is defined by two connected links with an angle each, which is visualized in Figure 2B. The first link is connected to a fixed position. In the swing up task, the goal of the policy is to bring the end point of the second link above a certain threshold height. The dynamics of the acrobot are defined as follows [sutton1995generalization]. Let and denote the angles of the first and second link, respectively. Their dynamics are given by
@@
In every experiment, is set to , to 1.5 and the initial states are sampled from . Observations of are wrapped between [-, ] and a trial terminates when or . Policies are discovered to solve the acrobot in three different experiments. In the first experiment the state of the acrobot is observable with noise and control force is only applied to the second link. In the second experiment the acrobot state is made partially observable, as the policies no longer observe the angular velocities. As in the first experiment, the control force is only applied to the second link. The third experiment tests for scalability of the policies under noisy observations. The policies apply control forces to both links in this experiment, therefore the policy should return two outputs. To increase the memory capacity of the dynamic policies, the latent size in dynamic policies is increased to three. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. As for the SHO, the parameters of the readout function are set to and in case of observation noise. In experiment with partial state observability, is again changed to exclude the unobserved variables and is shrunken to the correct dimension.
@@ 2.6.3 Continuous stirred tank reactor
The CSTR describes a chemical process following non-linear dynamics [uppal1974dynamic]. The CSTR consists of a main tank that contains a mixture of chemicals and a cooling jacket around the main tank. A visualization of the CSTR is presented in Figure 2C. To optimise the chemical process inside the main tank, the temperature has to be stabilised around a setpoint temperature. The temperature in the cooling jacket can be influenced directly by altering the input flow of coolant, which in turn affects the temperature in the main tank. The dynamics of the CSTR are defined as in [antonelli2003continuous], extended with a Wiener process to incorporate stochasticity. Let . The CSTR defined as a stochastic first-order system
@@
and parameters as defined in Appendix C.
@@
The policy has to minimise the distance between the temperature in the reactor and the setpoint temperature , while also minimising the control, i.e. the coolant flowrate. Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table 3 in Appendix C). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. The best dynamic symbolic policy evolved with GP is tested in a trial with changing setpoint temperatures, to analyse how the policy and latent state variables respond to these changes. To generate the observations, is defined such that the concentration is excluded from the observations and .
@@ 3 Results
We present a GP method for finding high-performing white-box symbolic policies for solving control problems. We consider both static and dynamic policies optimized by GP (GP-S and GP-D respectively), and test their performance by controlling three common control benchmark systems (as explained in detail in Section 2.6). We test whether GP evolves successful control policies for linear and non-linear environments, different observability levels, as well as changing environment conditions and multiple control dimensions. We compare the resulting control policies to those found from optimal control (where possible), random search (RS), and to a reference black-box model (NDEs).
@@ 3.1 Genetic programming evolves linear and non-linear symbolic policies
To study whether GP is able to efficiently evolve symbolic policies we first consider a stochastic harmonic oscillator (SHO) with a static environment setting. For this particular control problem the optimal LQG controller gives us a upper bound on the fitness (Figure 3A, black curve). Given the simplicity of the control problem, even random search can find control policies with close to optimal performance (Figure 3A, gray dots). However, these policies can take a long time and many experiments to find (Figure 3A, purple curve), resulting in low consistency of the validation scores across different runs. Ideally, evolution should find such policies faster and more consistently. Indeed, using evolution strategies to optimise NDEs allows us to quickly find control policies close to optimal (Figure 3A, green curve).
@@
To validate that GP can evolve symbolic policies that reach similar performance, we first optimized static symbolic policies (Figure 3A, blue curve). The initial average best fitness of GP-S is already close to the optimal fitness, which indicates that well-performing static policies already exist in the randomly sampled initial population, which is likely a result of the small search space of GP-S. However, from there on GP is not able to find policies reaching the optimal fitness, indicating that the forward policies do not have enough complexity to match the optimal controller. We next used GP to optimize dynamic symbolic policies (Figure 3A, orange curve). The average best fitness of GP-D begins significantly lower than GP-S does, reflecting the much larger search space and in turn lower probability to randomly sample a high-performing policy. However, from there on GP quickly converges to policies that are optimal or close to optimal, indicating that its internal dynamics are beneficial to deal with noisy observations. The GP-D solutions are also highly consistent across trials, although there was one outlier with less good performance (Figure 3A, blue dots).
The SHO can be stabilised with a linear controller, as can be deduced from the structure of the LQG controller. However, as most control problems require non-linear policies, it is important to demonstrate that such policies are also optimised with GP. Therefore, our method was applied to the acrobot swing-up task to demonstrate that non-linear policies are still evolved efficiently. Similarly to the previous experiment, random search performs much worse than the other methods in terms of average best fitness (Figure 3B, purple curve). The validation fitness is also worse, although there are some policies that overlap with the other methods (Figure 3B purple dots). NDE converges to a higher fitness than random search (Figure 3B, green curve), and the validation fitness of the best policies found with NDE is also better than any policy discovered with random search (Figure 3B, green dots).
Static and dynamic symbolic policies were evolved with GP to analyse whether similar performance as the NDE can be obtained. GP-S converges at a fitness value higher than random search, but lower than the NDE (Figure 3B, blue curve). The average best fitness of GP-D is lower than GP-S initially, but converges to a slightly higher fitness during evolution (Figure 3B, orange curve). The validation fitness of the best policies evolved with GP-D (Figure 3B, orange dots) and GP-S (Figure 3B, blue dots) overlap largely.
@@
Overall, the results in the experiment demonstrate that static and dynamic symbolic policies are effectively optimised with GP. Both high-performing linear and non-linear policies are evolved, therefore showing that GP is clearly advantageous for symbolic policy learning compared to random search. The black-box NDE is more consistent in finding high-performing policies, however the majority of the evolved symbolic policies are able to compete.
@@ 3.2 Dynamic symbolic policies remain robust under partial state observability
The previous experiments demonstrated that static and dynamic symbolic policies are successfully evolved for a linear and non-linear settings. However, these experiments assumed noisy observations of the full state, which is not representative for most real-life applications. We hypothesized that the memory capacity of the dynamic policies is particularly beneficial when the policies are confronted with partial observations of the state. To test this hypothesis, symbolic policies were evolved for stabilisation of the SHO and the acrobot swing-up under partial state observability, where the velocity is not observed in both environments.
@@
The SHO environment remains linear, therefore optimal LQG controller again represents the upper bound fitness value (Figure 3C, black curve). Random search was excluded in this experiment, because it could not consistently discover high-performing policies in the previous experiment (Figure 3A, purple curve). The NDE converges to a slightly lower fitness (Figure 3C, green curve) than it did in the full observable SHO (Figure 3A, green curve). However, the NDE still consistently produced policies that are close to optimal in terms of performance.
@@
As expected, the average best fitness of GP-S converges to a substantially poorer fitness than the NDE (Figure 3C, blue curve). The validation fitness of GP-S also shows that every policy obtains similar performance and fails to satisfy the task when the velocity is excluded from the observations (Figure 3C, blue dots). Differently from GP-S, well-performing dynamic symbolic policies are successfully optimised with GP in this experiment (Figure 3C, orange curve). The average best fitness of GP-D converges to a similar fitness as the NDE. Additionally, the validation fitness is consistent, and some policies even perform better than the NDE policies (Figure 3C, orange dots).
The dynamic symbolic policies perform significantly better than the static policies on the SHO with partial state observability. The acrobot swing-up experiment validates whether this difference in performance is also obtained in a non-linear environment. In this experiment, random search is again left out as baseline, because it could not produce comparable policies (Figure 3B, purple curve). The NDE approach converges to high-performing policies (Figure 3D, green curve), although at a slightly lower average best fitness than it did under full observability (Figure 3B, green curve). The validation fitness of the best NDE policies shows less consistency than in the previous experiments, as there are a few policies that perform substantially worse than the average (Figure 3D, green dots).
In the experiment with partial state observability in the SHO, the static policies failed to solve the task while the dynamic policies remained high-performing (Figure 3C). In the partially observable acrobot, similar findings are obtained. GP-S does not evolve a single policy that completes the swing up, as is observed from the constant fitness during evolution (Figure 3D, blue curve) and the lack of policies that obtain a validation fitness above the minimum fitness of 250 (Figure 3D, blue dots). The dynamic policies evolved with GP are able to solve the acrobot swing up (Figure 3D, orange curve). The validation fitness of GP-D shows consistency in performance among the discovered policies (Figure 3b, orange dots), even more so than the NDE baseline. The finding that dynamic symbolic policies are successful under partial observability was expected, because the best performing policy in the fully observable acrobot swing-up task did not include the velocity observations in its dynamics (Table 5, Acrobot, Observation noise).
@@
The limited performance of the static symbolic policies under partial state observability compared to the dynamic symbolic policies and NDE can be contributed to the lack of memory in the static policies. The memory capacity in the dynamic policies allows for estimation of the velocity, which turns out to be crucial to solve the SHO and acrobot tasks successfully. The memory capacity in the dynamic symbolic policy significantly increases the search space, but enables the policies to satisfy its goals in environments with partial observability.
@@ 3.3 Symbolic policies generalise to varying environments
In realistic applications, policies will often be confronted with partial observations. However, the environment of interest might also differ between trials or change over time, therefore the policies have to be able to generalise to different environment specifics. To validate that generalising symbolic policies can be evolved, the policies have to stabilise the SHO of which the parameters vary in every trial.
For this experiment, the upper bound fitness is again obtained with the optimal LQG controller, which is computed independently for every parameter setting (Figure 3E, black curve). The average best fitness of the NDE approach converges close to the upper bound fitness (Figure 3E, green curve). The validation fitness of the NDE policies shows that the evolved policies are high-performing in each evolutionary run (Figure 3E, green dots).
To validate whether symbolic policies can generalise to varying environment settings, static policies are first evolved (Figure 3E, blue curve). The average best fitness during evolution of GP-S does not converge near the upper bound fitness and the fitness obtained by the NDE. The validation fitness of the best static policies shows that a ceiling has been reached, as every policy is located close to the average (Figure 3E, blue dots). The average best fitness of GP-D converges to a similar fitness as the GP-S (Figure 3E, orange curve). However, the validation fitness of the best dynamic symbolic policies shows an interesting pattern (Figure 3E, orange dots). Many of the dynamic symbolic policies obtain a validation fitness that is better than any static policy, and some even compare with NDE policies. On the other hand, in some evolution processes of dynamic symbolic policies, GP failed to discover well-performing dynamic policies, as is observed from the validation fitness that is lower than any validation fitness obtained by GP-S.
The inconsistency of the evolution of dynamic symbolic policies can be explained by the large search space. However, when dynamic policies are evolved successfully, they show to perform better than static policies are capable of. The memory capacity allows dynamic policies to discover useful properties about the environment that increase the generalisability compared to static policies. In our results, there are more evolutionary runs that result in dynamic policies that exceed the performance of static policies than the other way around. Therefore the addition of the memory capacity in symbolic policies proves to be more beneficial than it is disadvantageous. Nonetheless, the black-box NDE consistently produces policies that outperform most symbolic policies.
@@ 3.4 Genetic programming evolves higher-dimensional policies
As the search space of symbolic policies is large, especially for the dynamic symbolic policies, it is important to validate whether high-performing are optimised efficiently when the control dimension increases. In this experiment, the policies can apply control forces to both links to accomplish the acrobot swing-up. The NDE converges to a slightly higher fitness during evolution than in the other experiments on the acrobot, because more force can be applied to solve the swing up (Figure 3F, green curve). The best NDE policies obtain consistent validation fitness (Figure 3F, green dots).
@@
Our method is first tested to produce static symbolic policies with two control outputs (Figure 3F, blue curve). The search space of static policies is still small compared to the dynamic policies, as is observed from the higher initial average best fitness compared to the other methods. The average best fitness of GP-S does not improve much during evolution, and converges at a worse fitness than the NDE. The validation fitness shows that most static policies perform similarly, but one evolved policy performs better than the other static policies (Figure 3F, blue dots). GP-D converges at a slightly better fitness than GP-S, but worse than NDE (Figure 3F, orange curve). The validation fitness obtained by dynamic symbolic policies mostly overlaps with that of the static symbolic policies and NDE policies, although a few policies are worse than any policy obtained by the other methods (Figure 3F, orange dots). This experiment demonstrates that GP still evolves higher-dimensional control policies that accomplish the acrobot swing up, even though the size of the search space increased even further.
3.5 Generalising symbolic policies are discovered for industrial application
In the final experiment, we apply our method to an industrial problem, the continuous stirred tank reactor (CSTR). A policy should stabilise the temperature in the reactor tank at a setpoint temperature, which is conducted by controlling the flow of coolant into the cooling jacket of the reactor. The goal of this experiment is to demonstrate that symbolic policies can be discovered that perform well in this realistic setting. The parameter setting in the CSTR is varied in each trial, therefore policies should generalise to different reactor specifics. GP is used to optimise static and dynamic policies again, and the policies are compared with NDE policies.

### v3

@@ 2.6.1 Stochastic harmonic oscillator
The SHO is a linear model that describes the position and velocity of an oscillating mass in a single dimension under random perturbations. The SHO is visualised in Fig. 3a and the dynamics of the SHO are given by
@@
with and . Here, punishes the policy for the distance between and some target state and penalises the control strength. The initial condition is randomly sampled from a normal distribution with zero mean and variance diag(3, 1) for the position and velocity respectively. The target position is chosen randomly by sampling uniformly between -3 and 3 in each trial, while the target velocity is always set to zero. As the optimal LQG control is linear, the function set includes basic arithmetic operators.
GP and the baselines are tested on the SHO in three different experiments. In the first experiment the policies interact with the standard setting of the SHO. In this experiment is specified with and and the state observations are noisy. In the second experiment the state of the SHO is partially observed, as the policies only observe the position while the velocity is unknown. remains the same in this experiment. In the third experiment, the policies are evaluated on generalisability. In each trial and are different, specifically and . In this experiment both the position and velocity are observed again. In the experiments with observation noise, the parameters of the readout function are set to and (see Equation 2). In experiments with partial state observability, is adapted such that only the relevant observations are returned and the dimension of is decreased to match the number of observations.
@@ 2.6.2 Acrobot swing up task
The acrobot is defined by two connected links, which is visualised in Fig. 3b. The first link is connected to a fixed position. In the swing up task, the goal of the policy is to bring the end point of the second link above a certain threshold height. The dynamics of the acrobot are defined as follows [sutton1995generalization]. Let and denote the angles of the first and second link, respectively. Their dynamics are given by
@@
In every experiment, is set to , to 1.5 and the initial states are sampled from . Observations of are wrapped between [-, ] and a trial terminates when or . Policies are discovered to solve the acrobot in three different experiments. In the first experiment the state of the acrobot is observable with noise and control force is only applied to the second link. In the second experiment the acrobot state is made partially observable, as the policies no longer observe the angular velocities. In the third experiment, the policies can apply independent forces to both links, and the angular velocities are observed again. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. As for the SHO, the parameters of the readout function are set to and in case of observation noise. In experiment with partial state observability, is again changed to exclude the unobserved variables and is shrunken to the correct dimension.
@@ 2.6.3 Continuous stirred tank reactor
The CSTR describes a chemical process following non-linear dynamics [uppal1974dynamic]. The CSTR consists of a main tank that contains a mixture of chemicals and a cooling jacket around the main tank. A visualisation of the CSTR is presented in Fig. 3c. To optimise the chemical process inside the main tank, the temperature has to be stabilised around a setpoint temperature. The temperature in the cooling jacket can be influenced directly by altering the input flow of coolant, which in turn affects the temperature in the main tank. The dynamics of the CSTR are defined as in [antonelli2003continuous], extended with a Wiener process to incorporate stochasticity. Let . The CSTR is defined as a stochastic first-order system
@@
and the parameters as defined in Table 4.
@@
The policy has to minimise the distance between the temperature in the reactor and the setpoint temperature , while also minimising the control, i.e. the coolant flowrate. Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table 4). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. To generate the observations, is defined such that the concentration is excluded from the observations and .
@@ 3 Results
We present a GP method for finding high-performing dynamic symbolic policies for solving control problems. The performance of a policy is tested by controlling three common control benchmark systems (as explained in detail in Section 2.6). We investigate whether GP evolves successful control policies for linear and non-linear environments, different observability levels, as well as changing environment conditions and multiple control dimensions. We compare the resulting dynamic symbolic policies to static symbolic policies, optimal control (where possible), random search (RS), and to a reference black-box model (NDEs).
@@ 3.1 Genetic programming evolves linear and non-linear symbolic policies
To study whether GP can successfully evolve dynamic symbolic policies, we first consider the SHO and acrobot environments with noisy full state observability (Fig. 4a,b). In the SHO, we consider the linear quadratic Gaussian (LQG) controller as the upper bound fitness (black curves). The fitness of random search improves during optimisation (purple curves), but does not converge near the upper bound nor the other methods. Random search does occasionally discover a policy that competes with the evolutionary approaches, but not consistently (purple dots). Moving forward, we will no longer consider random search as a baseline, but the results were similar across all tasks. NDEs converge to a high fitness quickly and consistently, approximating the upper bound fitness in the SHO (green curves and dots).
@@
To validate that symbolic policies reach similar performance as the NDE, we first optimised static symbolic policies (blue curves). The initial fitness of GP-S begins close to the upper bound fitness in the SHO, which indicates that well-performing static policies exist in the initial population, which is a result of the small search space of GP-S. However, from there, GP is not able to evolve policies reaching the optimal fitness, indicating that static policies are not complex enough to match the optimal controller. Similarly, in the acrobot swing-up task, GP-S begins at a high fitness, but does not improve much during evolution. In both tasks, GP-S converges significantly higher than RS, but is outperformed by the NDE in terms of fitness.
We next evolved dynamic symbolic policies (orange curves). In both tasks, the fitness of GP-D begins significantly lower than GP-S, due to the larger search space. However, GP-D quickly converges to a higher fitness than GP-S, approximating the upper bound fitness in the SHO. This indicates that the internal dynamics are able to correct for the noisy observations. Both the GP-S and GP-D policies are highly consistent across trials, with GP-D producing slightly better policies (blue and orange dots).
@@
Overall, these experiments demonstrates that static and dynamic symbolic policies are effectively optimised with GP, especially compared to random search. High-performing linear and non-linear policies are evolved that compete with the black-box NDE, although the NDE is more consistent.
@@ 3.2 Dynamic symbolic policies remain robust under partial state observability
The previous experiments demonstrated that static and dynamic symbolic policies are successfully evolved for linear and non-linear settings. However, these experiments assumed the full state is observed, which may not be representative for real-life applications. We hypothesised that the memory in dynamic policies is particularly beneficial when confronted with partial state observations. To test this hypothesis, symbolic policies were evolved for the SHO and acrobot environments under partial state observability, where the velocities are not observed (Fig. 4c,d).
@@
The SHO remains linear, therefore the LQG controller again represents the upper bound fitness (black curves). The NDE converges to a slightly lower fitness (green curves) compared to the experiments with full state observability. However, the NDE still consistently produced policies close to the upper bound in terms of performance, with a few policies that perform worse on the acrobot (green dots).
@@
As hypothesised, GP-S converges to a substantially lower fitness than under full state observability (blue curves). Every static symbolic policy fails to satisfy the task when the velocity is not observed (blue dots). Differently from GP-S, high-performing dynamic symbolic policies are still successfully evolved (orange curves). GP-D converges to a similar fitness as the NDE in both environments. Additionally, the validation fitness is consistent, and some policies even perform better than the NDE policies (orange dots).
@@
This experiment demonstrated that static symbolic policies perform poorly under partial state observability compared to the dynamic symbolic policies and NDE. This result can be contributed to the lack of memory in the static policies. Memory allows for estimation of the velocity, which enables policies to solve the SHO and acrobot tasks successfully. The latent state significantly increases the search space, but improves the robustness of the dynamic symbolic policies when confronted with partial observability.
@@ 3.3 Symbolic policies generalise to varying environments
Besides partial observations, realistic settings also require policies to generalise to varying and changing environments. To validate that generalising symbolic policies can be evolved, we test policies in the SHO with parameters that vary between trials (Fig. 4e). In this experiment, the upper bound fitness is again obtained with the LQG controller computed independently for each parameter setting (black curves). The NDE converges close to the upper bound fitness and is consistent across different runs (green curve and dots).
Static policies are first evolved to validate their generalisability (blue curve). GP-S does not converge near the upper bound fitness and the NDE. The validation fitness of the best static policies shows that a ceiling has been reached, as every policy is located close to the average (blue dots). GP-D converges to a similar fitness as GP-S (orange curve). However, the validation fitness of the dynamic symbolic policies shows an interesting pattern (orange dots). Many dynamic symbolic policies obtain a validation fitness that is better than any static policy, and some even compare with NDE policies. However, some dynamic symbolic policies obtain a validation fitness worse than any static policy.
The inconsistent evolution of dynamic symbolic policies is explained by the large search space. When dynamic policies are evolved successfully, they perform better than static policies. The memory allows dynamic policies to discover useful properties that increase generalisability. In our results, the majority of evolutionary runs resulted in dynamic policies that exceed the performance of static policies. Therefore the addition of the latent state in symbolic policies proves to be more beneficial than disadvantageous. Nonetheless, our approach does not consistently produce policies that perform comparably as the NDE.
@@ 3.4 Genetic programming evolves higher-dimensional policies
As the search space of symbolic policies is large, especially for the dynamic symbolic policies, it is important to validate whether high-performing are optimised efficiently when the control dimension increases. In this experiment, the policies can apply control forces to both links to accomplish the acrobot swing-up (Fig. 4f). The NDE converges to a slightly higher fitness during evolution than in the other experiments on the acrobot, because more force can be applied to solve the swing up (green curve). The best NDE policies obtain consistent validation fitness (green dots).
@@
Our method is first tested to produce static symbolic policies with two control outputs (blue curve). The search space of static policies is still small compared to the dynamic policies, as is observed from the higher initial average best fitness compared to the other methods. The average best fitness of GP-S does not improve much during evolution, and converges at a worse fitness than the NDE. The validation fitness shows that most static policies perform similarly, but one evolved policy performs better than the other static policies (blue dots). GP-D converges at a slightly better fitness than GP-S, but worse than NDE (orange curve). The validation fitness obtained by dynamic symbolic policies mostly overlaps with that of the static symbolic policies and NDE policies, although a few policies are worse than any policy obtained by the other methods (orange dots). This experiment demonstrates that GP still evolves higher-dimensional control policies that accomplish the acrobot swing up, even though the size of the search space was increased even further.
3.5 Generalising symbolic policies are discovered for an industrial application
In the final experiment, we apply our method to the continuous stirred tank reactor (CSTR), an industrial problem. The temperature in the reactor has to be stabilised at a setpoint temperature by controlling the flow of coolant into the cooling jacket (Fig. 4g). To increase the complexity of the problem, the environment setting is varied between trials and the policies receive partial state observations.